The sum of the ages of Fran and Laura is 33 years. In two years, 3 times Fran's age will be one year less than 4 times Laura's age. How old is each now?
Answer by TimothyLamb(4379)

F + L = 33
F = 33 - L
3(F + 2) = 4(L + 2) - 1
---
3F + 6 = 4L + 8 - 1
3F = 4L + 1
3(33 - L) = 4L + 1
99 - 3L = 4L + 1
7L = 98
Answer:
L = 14
F = 19
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php